OpenStudy (anonymous):

Segment AB has point A located at (5, 4). If the distance from A to B is 7 units, which of the following is the coordinate for point B? (5 points) Select one: a. (3, 2) b. (-2, 4) c. (2, 1) d. (12, 11)

OpenStudy (anonymous):

@MelissaHolmes @jim_thompson5910

OpenStudy (anonymous):

hi!

OpenStudy (anonymous):

helo can you help me on this one to

OpenStudy (anonymous):

ya, sure

OpenStudy (anonymous):

whoa... profile picture change there.... XD

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

d. because... (5, 4) 5 + 7 = 12 4 + 7 = 11 thus... (12, 11)

OpenStudy (anonymous):

@iambatman do agree

jimthompson5910 (jim_thompson5910):

Hint: use the distance formula \[\Large d = \sqrt{\left(x_{2}-x_{1}\right)^2+\left(y_{2}-y_{1}\right)^2}\]

OpenStudy (anonymous):

soooo.... was i right? i never learned anything about a formula....

jimthompson5910 (jim_thompson5910):

(x1,y1) = (5,4) (x2,y2) = (x,y) d = 7 so... \[\Large d = \sqrt{\left(x_{2}-x_{1}\right)^2+\left(y_{2}-y_{1}\right)^2}\] \[\Large 7 = \sqrt{\left(x-5\right)^2+\left(y-4\right)^2}\]

OpenStudy (anonymous):

so you would multiply out the squared stuff and solve for x and y?

jimthompson5910 (jim_thompson5910):

Now let's see if (12,11) is 7 units away from point A. Plug in (x,y) = (12,11) \[\Large 7 = \sqrt{\left(x-5\right)^2+\left(y-4\right)^2}\] \[\Large 7 = \sqrt{\left(12-5\right)^2+\left(11-4\right)^2}\] \[\Large 7 = \sqrt{\left(7\right)^2+\left(7\right)^2}\] \[\Large 7 = \sqrt{49+49}\] \[\Large 7 = \sqrt{98}\] \[\Large 7 = 9.89949493\] So because that last equation is false, this means that (12,11) is NOT 7 units away from point A

OpenStudy (anonymous):

hmm... so how would you find out what the actual answer is?

jimthompson5910 (jim_thompson5910):

plug in each point and see which satisfy the equation \[\Large 7 = \sqrt{\left(x-5\right)^2+\left(y-4\right)^2}\]

OpenStudy (anonymous):

oh.... so no way to actually solve?

OpenStudy (anonymous):

what should i go with

jimthompson5910 (jim_thompson5910):

unfortunately no because there are infinitely many points that are 7 units away from A

jimthompson5910 (jim_thompson5910):

if we could replace y with some expression in terms of x, then we can solve

OpenStudy (anonymous):

hmmm.... i see..... ok

OpenStudy (anonymous):

ok then ill get to the next one then ok

OpenStudy (anonymous):

so we each do one?

OpenStudy (anonymous):

ok im about to post it here

OpenStudy (anonymous):

Two boats start their journey from the same point A and travel along directions AC and AD, as shown below.

OpenStudy (anonymous):

What is the distance, CD, between the boats? (5 points) Select one: a. 461.9 ft b. 530.9 ft c. 646.4 ft d. 325.5 ft

OpenStudy (anonymous):

@jim_thompson5910 do u know this?

jimthompson5910 (jim_thompson5910):

yes, it turns out that B is the answer because... \[\Large 7 = \sqrt{\left(x-5\right)^2+\left(y-4\right)^2}\] \[\Large 7 = \sqrt{\left(-2-5\right)^2+\left(4-4\right)^2}\] \[\Large 7 = \sqrt{\left(-7\right)^2+\left(0\right)^2}\] \[\Large 7 = \sqrt{49+0}\] \[\Large 7 = \sqrt{49}\] \[\Large 7 = 7\] So (-2,4) is one point that is 7 units away from A(5,4)

OpenStudy (anonymous):

is 5,4 the answer

OpenStudy (anonymous):

hmm... ok :D

OpenStudy (anonymous):

ok what about this one

OpenStudy (anonymous):

good job @jim_thompson5910 you are very knowledgeable :D

OpenStudy (anonymous):

i just posted

jimthompson5910 (jim_thompson5910):

hint: first find the lengths of BC and BD to get CD, you subtract, so CD = BD - BC

jimthompson5910 (jim_thompson5910):

you will use the idea that \[\Large \tan(\theta) = \frac{\text{opp}}{\text{adj}}\]

OpenStudy (anonymous):

so what would it be @jim_thompson5910

jimthompson5910 (jim_thompson5910):

are you able to find BC or BD?

OpenStudy (anonymous):

NO i was thinking it was A but maybe im wrong

jimthompson5910 (jim_thompson5910):

I'll show you how to find BC \[\Large \tan(\theta) = \frac{\text{opp}}{\text{adj}}\] \[\Large \tan(60) = \frac{400}{BC}\] \[\Large BC\tan(60) = 400\] \[\Large BC = \frac{400}{\tan(60)}\]

OpenStudy (anonymous):

i have three more questions after thi one ok

OpenStudy (anonymous):

actually four hope you guys can help you really have been good to me

OpenStudy (anonymous):

1, 249.842 is what i got. but that doesn't seem right... @jim_thompson5910 ?

OpenStudy (anonymous):

oh, and sure @prowrestler i'll help the best i can :D

jimthompson5910 (jim_thompson5910):

Finding BD \[\Large \tan(\theta) = \frac{\text{opp}}{\text{adj}}\] \[\Large \tan(30) = \frac{400}{BD}\] \[\Large BD\tan(30) = 400\] \[\Large BD = \frac{400}{\tan(30)}\]

jimthompson5910 (jim_thompson5910):

Now subtract \[\Large CD = BD - BC\] \[\Large CD = \frac{400}{\tan(30)} - \frac{400}{\tan(60)}\] \[\Large CD \approx 461.8802153517 ... \ \text{Use a calculator here}\]

OpenStudy (anonymous):

So the answer is A

jimthompson5910 (jim_thompson5910):

for that last line, you type in "400/tan(30) - 400/tan(60)" without quotes

jimthompson5910 (jim_thompson5910):

If you do not have a calculator, google is a great calculator Just type in "400/tan(30 degrees) - 400/tan(60 degrees)" without quotes into google to get https://www.google.com/search?q=400/tan(30%20degrees)%20-%20400/tan(60%20degrees)

OpenStudy (anonymous):

my calculator came out with -1, 312.290?????

jimthompson5910 (jim_thompson5910):

you need to be in degree mode MelissaHolmes

OpenStudy (anonymous):

aahhhh... ok i see.... that makes sense.....

OpenStudy (anonymous):

So the answer is A right

jimthompson5910 (jim_thompson5910):

and google by default is in radian mode, so you have to tack on the word "degrees" in each tangent to force degree mode in google

jimthompson5910 (jim_thompson5910):

yes prowrestler

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

couple more guys and gals

OpenStudy (anonymous):

ok!

OpenStudy (anonymous):

Find the perimeter of a quadrilateral with vertices at C (-2, 1), D (2, 4), E (5, 0), and F (1, -3). Round your answer to the nearest hundredth when necessary. (5 points) Select one: a. 12 units b. 16 units c. 20 units d. 24 units

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

ummm.... this is completely new to me.... sorry...... :( maybe post it as a new question and see if anyone else can help u...

OpenStudy (anonymous):

I want to see if @jim_thompson5910 can help hes smart

OpenStudy (anonymous):

and you are to:)

OpenStudy (anonymous):

i know, he IS smart... he's helping someone else currently apparently

OpenStudy (anonymous):

oh, and thanks!

OpenStudy (anonymous):

stay here though please

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

Find the perimeter of a quadrilateral with vertices at C (-2, 1), D (2, 4), E (5, 0), and F (1, -3). Round your answer to the nearest hundredth when necessary. (5 points) Select one: a. 12 units b. 16 units c. 20 units d. 24 units

OpenStudy (anonymous):

oh, u know what... i'll graph the points and try to figure it out and be helpful :D

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i love you for this

OpenStudy (anonymous):

@jim_thompson5910 can you help now:)

OpenStudy (anonymous):

ok so i graphed this and i got this:

OpenStudy (anonymous):

hey fan me back and give me a medal pleeease:)

OpenStudy (anonymous):

so maybe around 24 units?

OpenStudy (anonymous):

sure!

OpenStudy (anonymous):

can i have a medal to

OpenStudy (anonymous):

medal?

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

what do you look like

OpenStudy (anonymous):

@jim_thompson5910 heres the question

OpenStudy (anonymous):

Find the perimeter of a quadrilateral with vertices at C (-2, 1), D (2, 4), E (5, 0), and F (1, -3). Round your answer to the nearest hundredth when necessary. (5 points) Select one: a. 12 units b. 16 units c. 20 units d. 24 units

jimthompson5910 (jim_thompson5910):

MelissaHolmes it looks like you used geogebra to plot those points. You can use geogebra to calculate the distances/lengths and find the perimeter.

OpenStudy (anonymous):

we thought it was 24 units

OpenStudy (anonymous):

i can? and yeah... i did use geogebra...

jimthompson5910 (jim_thompson5910):

see attached

OpenStudy (anonymous):

whoa.... okay... i guess you can.....

OpenStudy (anonymous):

A segment with endpoints A (3, 4) and C (5, 11) is partitioned by a point B such that AB and BC form a 2:3 ratio. Find B. (5 points) Select one: a. (3.8, 6.8) b. (3.9, 4.8) c. (4.2, 5.6) d. (4.3, 5.9)

jimthompson5910 (jim_thompson5910):

it's possible you just didn't have the "algebra" window open to see the segment lengths

OpenStudy (anonymous):

ok.... maybe...

OpenStudy (anonymous):

oh, yeah that makes sense :D

OpenStudy (anonymous):

@jim_thompson5910 can u help?

jimthompson5910 (jim_thompson5910):

A (3, 4) and C (5, 11) have x coordinates of 3 and 5 respectively. That x difference is 5-3 = 2 units The given ratio is 2:3 which means we'll have a total of 2+3 = 5 parts Split that 2 unit gap into 5 parts to get 2/5 = 0.4 ------------------------------------------------------ The y difference is 11 - 4 = 7 Split that gap into 5 pieces to get 7/5 = 1.4 ------------------------------------------------------ Because the first number of the ratio is 2, we will move 2 increments in both the x and y directions (starting from point A) so this means x = 3 ----> 3 + 2*(0.4) = 3.8 y = 4 ----> 4 + 2*(1.4) = 6.8 meaning that point B is at (3.8, 6.8) See attached to have a look at what I'm describing

OpenStudy (anonymous):

so the answer is A right